# Reasoning Under Threat: Symbolic and Neural Techniques for Cybersecurity Verification

### v1

@@ 1 Introduction
Cybersecurity has become a paramount concern in our increasingly interconnected world, with diverse application domains such as healthcare, supply chain management, and the Internet of Things facing sophisticated and evolving threats [66, 40, 24]. The complexity and security-critical nature of modern cyber systems necessitate rigorous analysis and verification to ensure their correctness, reliability, and resilience against potential vulnerabilities [41, 23]. Logic and automated reasoning play a pivotal role in providing formal guarantees for these crucial properties of cybersecurity [22, 19, 30]. By employing various logical frameworks and automated techniques, researchers and practitioners can formally model, specify, and verify different aspects of cyber systems, ranging from the intricate logic of security protocols to the correctness of access control mechanisms and the robustness against attacks [17, 34, 2]. This survey aims to provide a comprehensive overview of the current state of research at the intersection of logic and automated reasoning for cybersecurity. It delves into the diverse types of logic utilized, the automated reasoning techniques employed, the specific applications addressed, the limitations encountered, and the promising directions for future research in this rapidly evolving field [20, 49, 48].
@@
This survey advances the state of research at the nexus of formal logic, automated reasoning, and cybersecurity by offering a comprehensive and technically grounded synthesis of existing literature. Through an in-depth analysis of seminal works and recent advancements [68, 12, 51], the paper unpacks how logical systems and reasoning tools have been harnessed to formalize, analyze, and secure complex cyber infrastructures. Beyond mapping current capabilities, it surfaces critical blind spots in scalability, tool integration, and applicability to emerging domains [73, 11]. The survey further charts a forward trajectory by outlining technically rigorous and strategically relevant research directions, including hybrid AI-logical methods and compositional security verification, grounded in formal and mathematical principles [72, 71]. In doing so, it provides both a snapshot of the field’s current maturity and a blueprint for its evolution.
@@ 2 A Unified View of Logical Frameworks and Reasoning Tools in Cybersecurity
Formal logic and automated reasoning serve as the cornerstones of rigorous cybersecurity analysis. Logic provides the formal languages needed to specify security properties such as confidentiality, authentication, and integrity, while automated reasoning techniques enable the systematic verification of these properties at scale [32, 25]. Different logical systems offer varying expressive power and suitability for particular classes of security challenges [26, 39].
@@
First-order logic (FOL) has long served as a foundation for axiomatic modeling of security properties and system behavior [57, 1]. Temporal logics, including Linear Temporal Logic (LTL), are widely used for specifying and verifying time-dependent security properties, such as protocol state transitions and replay attack resilience [47, 18, 3, 2]. Deontic logic, which expresses permissions, obligations, and prohibitions, is especially relevant for access control policies, compliance verification, and security governance [62, 42]. More expressive frameworks, such as higher-order logic (HOL), enable formalization of complex system architectures and layered security models [45].
@@
These logical systems are realized through practical formal methods and reasoning tools that automate verification tasks. The K framework, Lean-based Clear, and Coq-based ConCert enable rigorous specification and formal verification of security protocols and software modules [30, 8, 41]. Advanced efforts such as those by Xu et al. [68] and Cai et al. [12] demonstrate how refinement calculus and separation logic can be enhanced with automated tools for scalable, sound verification. These efforts build on a rich tradition of foundational verification frameworks [4, 7, 58].
@@
Automated reasoning also plays a key role in discharge of proof obligations and satisfiability checks. SMT solvers such as Z3 are tightly integrated into security verification pipelines in tools like Lean and K [69, 50]. Recent work explores the use of SMT-based optimization techniques in static analysis for vulnerability detection and protocol hardening [56].
@@
Theorem proving and static analysis have been employed to verify both low-level code and high-level security policies [64, 20]. Applications of formal reasoning span from graph-based encrypted traffic analysis [51] to secure communication protocol verification using vector modulation logics [59]. Cyber-physical and federated systems—including IoT architectures—pose unique security challenges that have been addressed through logic-based verification frameworks emphasizing explainability and privacy [53, 37, 43, 48].
@@
The interaction between AI and formal logic opens new avenues for specification inference, vulnerability detection, and proof automation. Hybrid neural-symbolic approaches are gaining traction for security applications [39, 11], including recent efforts that explore LLM-guided proof synthesis and formal explanation generation [72].
@@
Security in distributed systems, particularly secure communication protocols and consensus mechanisms, has also benefited from formal specification and model checking. Tools such as PRISM have been used to evaluate probabilistic security guarantees under partial synchrony or Byzantine assumptions [55, 60]. Formal analysis of incentive compatibility, game-theoretic adversarial strategies, and trust assumptions is exemplified by the CheckMate framework and related work [63, 70].
@@
Finally, the application of formal verification in industry—such as in high-assurance blockchain platforms like Cardano and Tezos—demonstrates the growing importance of logic-based methodologies for real-world secure systems. Mi-Cho-Coq’s verification of the Michelson language [28, 13] and IOHK’s Haskell-based research [65, 52] reflect a broader trend toward adopting formal methods as first-class security engineering tools.
@@ 3.1 Verified Security Protocols: From Specification to Implementation
Security protocols must guarantee properties such as confidentiality, integrity, and authentication even under adversarial conditions. Formal verification plays a pivotal role in ensuring these guarantees through precise, logic-based specifications and automated proof frameworks [73, 15]. Environments such as Coq (via ConCert), Lean (via Clear), and the K framework offer powerful platforms for protocol verification [30, 69, 8]. These tools allow protocols to be encoded as logical assertions, supporting the verification of key properties through theorem proving and symbolic automation.
The use of refinement types has further increased the expressiveness and safety of formal specifications [25, 29]. Security-focused verification techniques often incorporate SMT solvers and model checkers to validate complex properties such as key secrecy, mutual authentication, and tamper resistance [45, 31]. Approaches that combine symbolic execution, static analysis, and logic-based verification help reduce false positives and negatives, improving assurance and automation [44, 26, 21].
Scaling these techniques to real-world protocols requires modularity and compositionality. Work by Bride et al. [35, 16] and Hou et al. [33] demonstrates frameworks for compositional verification of large-scale systems. Translation from verified specifications to executable code is a critical step in bridging theory and deployment, supported by tools such as those in [27]. Verified implementations have been deployed in sectors including finance and healthcare, as shown in [9, 5], and verification has extended to cover real-time and resource-constrained environments such as embedded and IoT systems [14].
@@ 3.2 Logical Modeling and Analysis of Secure Communication Protocols
Designing secure communication protocols requires formal models that accurately represent state transitions, adversarial capabilities, and probabilistic behavior. Logical methods—including state machines, process algebras, temporal logic, and epistemic models—enable rigorous design and analysis of such systems. Probabilistic frameworks like Continuous Time Markov Chains (CTMCs) support the modeling of stochastic security behavior and uncertain message delivery, as demonstrated in the analysis of randomized protocols [6].
Temporal logics such as LTL and CTL are particularly effective for expressing critical properties like safety and liveness in secure communication scenarios [67, 18]. Model checking techniques have been applied extensively to analyze Byzantine Fault-Tolerant (BFT) protocols and ensure robustness in the presence of faulty or malicious participants [54, 23]. Probabilistic model checking tools such as PRISM and PRISM+ have been used to explore network-level properties and failure scenarios [70].
More expressive models such as threshold automata and temporal epistemic logic enable reasoning about distributed trust, partial knowledge, and cryptographic guarantees [63, 57]. As secure communication protocols grow in complexity, compositional reasoning becomes vital. Logic-based frameworks that support asynchronous communication and hybrid systems allow modeling of precise timing and ordering constraints in secure messaging [61].
The growing prominence of federated and cross-domain systems introduces new challenges for security reasoning. Federated learning and multi-party protocols must coordinate trust, privacy, and communication under incomplete knowledge. Research on privacy-preserving coordination and federated trust models [38, 50] is increasingly important. Game-theoretic reasoning, combined with SAT/SMT-based symbolic utilities, has been applied to analyze strategic behavior and incentive structures in adversarial communication settings [20, 48, 52].
@@ 3.3 Automated Reasoning for Vulnerability Detection and Threat Analysis
Identifying vulnerabilities in complex systems requires systematic analysis of control flow, data dependencies, and threat surfaces. Automated reasoning techniques—especially those built on symbolic logic—enable scalable vulnerability detection in software, hardware, and network infrastructures. Tools like Mythril and Oyente apply symbolic execution to enumerate execution paths and detect vulnerabilities such as reentrancy, overflow, and unauthorized access [10, 39]. These systems use SMT solvers and constraint encodings to reason about safety violations under adversarial inputs.
Recent approaches expand these capabilities through modular engines, theorem proving integration, and semantic reasoning backends, improving precision and coverage [73, 72]. Static analysis methods based on logical inference have been used to verify security properties such as non-interference and secure information flow [64, 20]. Graph-based reasoning over encrypted communication flows has proven effective for network vulnerability assessment [51], while formal logics like vector modulation have enabled analysis of low-level communication channels [59].
Threat modeling, a key step in proactive security engineering, is often grounded in logic-based policy languages and epistemic frameworks. These approaches have been used to formally specify access control [56], analyze resilience to denial-of-service attacks [46], and synthesize defense strategies through logic programming [49]. Integration of fuzzing, bounded exploration, and Markovian models further enhances coverage of potential attack vectors [65, 39].
An emerging direction combines automated reasoning with machine learning to detect complex vulnerabilities and support intelligent security monitoring. AI-enhanced vulnerability detection has shown promise in identifying subtle flaws in large codebases, especially when integrated with logical verification engines [1, 41]. Tools that leverage formal logic alongside neural models can improve explainability and reduce verification overhead [36, 45]. Hybrid learning-verification pipelines, as explored in [71], represent a promising research trajectory. In distributed settings such as federated systems, reasoning about vulnerabilities requires logic-based analysis of confidentiality, integrity, and trust relations across multiple entities [43, 37, 47].
@@ 4 Toward Unified Reasoning for Cybersecurity: Gaps, Challenges, and Underexplored Directions
Despite significant advances in formal methods, automated reasoning, and cybersecurity engineering, there remains a notable fragmentation in the literature addressing their intersection. While numerous studies explore individual components—such as vulnerability detection, security protocol verification, or logic-based policy enforcement—comprehensive surveys that synthesize the role of formal logic and automated reasoning across cybersecurity domains are scarce. Existing surveys often prioritize system performance [73], smart contract security [11], or symbolic analysis [41], but few address the integration of logical frameworks and reasoning automation as core pillars of cybersecurity design and assurance.
@@
A critical gap lies in the limited adoption of expressive and modular logical systems in security analysis. While propositional and first-order logic are widely used, more expressive formalisms—such as temporal epistemic logic, probabilistic deontic logic, and hybrid modal logics—remain largely underutilized, despite their potential to model knowledge, uncertainty, obligation, and time in adversarial settings [25, 29]. Temporal epistemic logic, for example, can naturally capture information flow, authentication handshakes, and adversary observability across protocol steps. Probabilistic deontic frameworks can encode uncertainty in system obligations (e.g., rate-limited access control), yet these logics are seldom integrated into real-world security tools.
@@
Another underexplored area is the development of unified verification pipelines that combine diverse reasoning strategies. While tools for symbolic execution, SMT solving, model checking, and theorem proving exist in isolation, few frameworks integrate these paradigms into cohesive architectures for holistic cybersecurity assurance. Hybrid static-dynamic verification frameworks, compositional reasoning environments, and multi-layered specification languages remain topics of active research but lack industrial adoption and tooling maturity. Recent efforts to bridge this gap, such as the integration of symbolic execution engines with refinement-type systems or proof assistants [26, 30], point in promising directions but require further generalization and scalability engineering.
@@
The application of formal methods in emerging domains—such as IoT security, federated systems, and cyber-physical infrastructures—also suffers from a lack of comprehensive reasoning frameworks. Much of the current work is focused on traditional computing environments, with limited coverage of context-aware or resource-constrained systems [57, 24]. These domains introduce unique security challenges such as real-time constraints, decentralized trust assumptions, and adversarial control over physical interfaces. Although logic-based analysis has been extended to federated models [37, 43] and constrained networks [14], there is a need for logic-aware formal DSLs and compositional specification languages tailored to these environments.
@@
Scalability remains a fundamental bottleneck. While symbolic reasoning tools like SMT solvers and model checkers offer automation, their application to large-scale, real-world systems is often limited by computational overhead and manual specification burdens. Verification frameworks struggle to keep up with the dynamic, heterogeneous nature of modern cyber-infrastructures. Compositional verification—using techniques such as assume-guarantee reasoning, modular contracts, and indexed logical types—offers a potential pathway forward [48, 35]. However, existing tools only partially support these methods, and integration into agile development pipelines remains rare.
Emerging research at the intersection of AI and formal methods offers a compelling vision for next-generation reasoning systems, but it is still in its infancy. Neural-symbolic methods for specification inference, counterexample generation, and proof synthesis show promise [72, 71], yet current prototypes lack formal correctness guarantees or clear trust boundaries between learning components and deductive backends. Trust-aware pipelines, such as those based on proof-carrying code or property-carrying predictions, could help mitigate these concerns, but further theoretical development is needed.

### v2

@@ 1 Introduction
Cybersecurity has become a paramount concern in our increasingly interconnected world, with diverse application domains such as healthcare, supply chain management, and the Internet of Things [1] facing sophisticated and evolving threats. The complexity and security-critical nature of modern cyber systems necessitate rigorous analysis and verification to ensure their correctness, reliability, and resilience against potential vulnerabilities. Logic and automated reasoning play a pivotal role in providing formal guarantees for these crucial properties of cybersecurity. By employing various logical frameworks and automated techniques, researchers and practitioners can formally model, specify, and verify different aspects of cyber systems, ranging from the intricate logic of security protocols to the correctness of access control mechanisms and the robustness against attacks. This survey aims to provide a comprehensive overview of the current state of research at the intersection of logic and automated reasoning for cybersecurity. It delves into the diverse types of logic utilized, the automated reasoning techniques employed, the specific applications addressed, the limitations encountered, and the promising directions for future research in this rapidly evolving field.
@@
This survey advances the state of research at the nexus of formal logic, automated reasoning, and cybersecurity by offering a comprehensive and technically grounded synthesis of existing literature. Through an in-depth analysis of seminal works and recent advancements, the paper unpacks how logical systems and reasoning tools have been harnessed to formalize, analyze, and secure complex cyber infrastructures. Beyond mapping current capabilities, it surfaces critical blind spots in scalability, tool integration, and applicability to emerging domains. The survey further charts a forward trajectory by outlining technically rigorous and strategically relevant research directions, including hybrid AI-logical methods and compositional security verification, grounded in formal and mathematical principles. In doing so, it provides both a snapshot of the field’s current maturity and a blueprint for its evolution.
@@ 2 A Unified View of Logical Frameworks and Reasoning Tools in Cybersecurity
Formal logic and automated reasoning serve as the cornerstones of rigorous cybersecurity analysis. Logic provides the formal languages needed to specify security properties such as confidentiality, authentication, and integrity, while automated reasoning techniques enable the systematic verification of these properties at scale. Different logical systems offer varying expressive power and suitability for particular classes of security challenges.
@@
First-order logic (FOL) has long served as a foundation for axiomatic modeling of security properties and system behavior. Temporal logics, including Linear Temporal Logic (LTL), are widely used for specifying and verifying time-dependent security properties, such as protocol state transitions and replay attack resilience. Deontic logic, which expresses permissions, obligations, and prohibitions, is especially relevant for access control policies, compliance verification, and security governance. More expressive frameworks, such as higher-order logic (HOL), enable formalization of complex system architectures and layered security models.
@@
These logical systems are realized through practical formal methods and reasoning tools that automate verification tasks. The K framework [3], Lean-based Clear, and Coq-based ConCert enable rigorous specification and formal verification of security protocols and software modules. Advanced efforts such as those by Cai et al. demonstrate how refinement calculus and separation logic can be enhanced with automated tools for scalable, sound verification. These efforts build on a rich tradition of foundational verification frameworks.
@@
Automated reasoning also plays a key role in discharge of proof obligations and satisfiability checks. SMT solvers such as Z3 [2] are tightly integrated into security verification pipelines in tools like Lean and K. Recent work explores the use of SMT-based optimization techniques in static analysis for vulnerability detection and protocol hardening.
@@
Theorem proving and static analysis have been employed to verify both low-level code and high-level security policies. Applications of formal reasoning span from graph-based encrypted traffic analysis to secure communication protocol verification using vector modulation logics. Cyber-physical and federated systems—including IoT architectures—pose unique security challenges that have been addressed through logic-based verification frameworks emphasizing explainability and privacy.
@@
The interaction between AI and formal logic opens new avenues for specification inference, vulnerability detection, and proof automation. Hybrid neural-symbolic approaches are gaining traction for security applications, including recent efforts that explore LLM-guided proof synthesis and formal explanation generation.
@@
Security in distributed systems, particularly secure communication protocols and consensus mechanisms, has also benefited from formal specification and model checking. Tools such as PRISM have been used to evaluate probabilistic security guarantees under partial synchrony or Byzantine assumptions. Formal analysis of incentive compatibility, game-theoretic adversarial strategies, and trust assumptions is exemplified by the CheckMate framework and related work.
@@
Finally, the application of formal verification in industry—such as in high-assurance blockchain platforms like Cardano and Tezos—demonstrates the growing importance of logic-based methodologies for real-world secure systems. Mi-Cho-Coq’s verification of the Michelson language and IOHK’s Haskell-based research reflect a broader trend toward adopting formal methods as first-class security engineering tools.
@@ 3.1 Verified Security Protocols: From Specification to Implementation
Security protocols must guarantee properties such as confidentiality, integrity, and authentication even under adversarial conditions. Formal verification plays a pivotal role in ensuring these guarantees through precise, logic-based specifications and automated proof frameworks. Environments such as Coq (via ConCert), Lean (via Clear), and the K framework offer powerful platforms for protocol verification. These tools allow protocols to be encoded as logical assertions, supporting the verification of key properties through theorem proving and symbolic automation.
The use of refinement types has further increased the expressiveness and safety of formal specifications. Security-focused verification techniques often incorporate SMT solvers and model checkers to validate complex properties such as key secrecy, mutual authentication, and tamper resistance. Approaches that combine symbolic execution, static analysis, and logic-based verification help reduce false positives and negatives, improving assurance and automation.
Scaling these techniques to real-world protocols requires modularity and compositionality. The literature demonstrates frameworks for compositional verification of large-scale systems. Translation from verified specifications to executable code is a critical step in bridging theory and deployment, supported by tools such as those in. Verified implementations have been deployed in sectors including finance and healthcare, as shown in, and verification has extended to cover real-time and resource-constrained environments such as embedded and IoT systems.
@@ 3.2 Logical Modeling and Analysis of Secure Communication Protocols
Designing secure communication protocols requires formal models that accurately represent state transitions, adversarial capabilities, and probabilistic behavior. Logical methods—including state machines, process algebras, temporal logic, and epistemic models—enable rigorous design and analysis of such systems. Probabilistic frameworks like Continuous Time Markov Chains (CTMCs) support the modeling of stochastic security behavior and uncertain message delivery, as demonstrated in the analysis of randomized protocols.
Temporal logics such as LTL and CTL are particularly effective for expressing critical properties like safety and liveness in secure communication scenarios. Model checking techniques have been applied extensively to analyze Byzantine Fault-Tolerant (BFT) protocols and ensure robustness in the presence of faulty or malicious participants. Probabilistic model checking tools such as PRISM and PRISM+ have been used to explore network-level properties and failure scenarios.
More expressive models such as threshold automata and temporal epistemic logic enable reasoning about distributed trust, partial knowledge, and cryptographic guarantees. As secure communication protocols grow in complexity, compositional reasoning becomes vital. Logic-based frameworks that support asynchronous communication and hybrid systems allow modeling of precise timing and ordering constraints in secure messaging.
The growing prominence of federated and cross-domain systems introduces new challenges for security reasoning. Federated learning and multi-party protocols must coordinate trust, privacy, and communication under incomplete knowledge. Research on privacy-preserving coordination and federated trust models is increasingly important. Game-theoretic reasoning, combined with SAT/SMT-based symbolic utilities, has been applied to analyze strategic behavior and incentive structures in adversarial communication settings.
@@ 3.3 Automated Reasoning for Vulnerability Detection and Threat Analysis
Identifying vulnerabilities in complex systems requires systematic analysis of control flow, data dependencies, and threat surfaces. Automated reasoning techniques—especially those built on symbolic logic—enable scalable vulnerability detection in software, hardware, and network infrastructures. Tools like Mythril and Oyente apply symbolic execution to enumerate execution paths and detect vulnerabilities such as reentrancy, overflow, and unauthorized access. These systems use SMT solvers and constraint encodings to reason about safety violations under adversarial inputs.
Recent approaches expand these capabilities through modular engines, theorem proving integration, and semantic reasoning backends, improving precision and coverage. Static analysis methods based on logical inference have been used to verify security properties such as non-interference and secure information flow. Graph-based reasoning over encrypted communication flows has proven effective for network vulnerability assessment, while formal logics like vector modulation have enabled analysis of low-level communication channels.
Threat modeling, a key step in proactive security engineering, is often grounded in logic-based policy languages and epistemic frameworks. These approaches have been used to formally specify access control, analyze resilience to denial-of-service attacks, and synthesize defense strategies through logic programming. Integration of fuzzing, bounded exploration, and Markovian models further enhances coverage of potential attack vectors.
An emerging direction combines automated reasoning with machine learning to detect complex vulnerabilities and support intelligent security monitoring. AI-enhanced vulnerability detection has shown promise in identifying subtle flaws in large codebases, especially when integrated with logical verification engines. Tools that leverage formal logic alongside neural models can improve explainability and reduce verification overhead. Hybrid learning-verification pipelines, as explored in, represent a promising research trajectory. In distributed settings such as federated systems, reasoning about vulnerabilities requires logic-based analysis of confidentiality, integrity, and trust relations across multiple entities.
@@ 4 Toward Unified Reasoning for Cybersecurity: Gaps, Challenges, and Underexplored Directions
Despite significant advances in formal methods, automated reasoning, and cybersecurity engineering, there remains a notable fragmentation in the literature addressing their intersection. While numerous studies explore individual components—such as vulnerability detection, security protocol verification, or logic-based policy enforcement—comprehensive surveys that synthesize the role of formal logic and automated reasoning across cybersecurity domains are scarce. Existing surveys often prioritize system performance, smart contract security, or symbolic analysis, but few address the integration of logical frameworks and reasoning automation as core pillars of cybersecurity design and assurance.
@@
A critical gap lies in the limited adoption of expressive and modular logical systems in security analysis. While propositional and first-order logic are widely used, more expressive formalisms—such as temporal epistemic logic, probabilistic deontic logic, and hybrid modal logics—remain largely underutilized, despite their potential to model knowledge, uncertainty, obligation, and time in adversarial settings. Temporal epistemic logic, for example, can naturally capture information flow, authentication handshakes, and adversary observability across protocol steps. Probabilistic deontic frameworks can encode uncertainty in system obligations (e.g., rate-limited access control), yet these logics are seldom integrated into real-world security tools.
@@
Another underexplored area is the development of unified verification pipelines that combine diverse reasoning strategies. While tools for symbolic execution, SMT solving, model checking, and theorem proving exist in isolation, few frameworks integrate these paradigms into cohesive architectures for holistic cybersecurity assurance. Hybrid static-dynamic verification frameworks, compositional reasoning environments, and multi-layered specification languages remain topics of active research but lack industrial adoption and tooling maturity. Recent efforts to bridge this gap, such as the integration of symbolic execution engines with refinement-type systems or proof assistants, point in promising directions but require further generalization and scalability engineering.
@@
The application of formal methods in emerging domains—such as IoT security, federated systems, and cyber-physical infrastructures—also suffers from a lack of comprehensive reasoning frameworks. Much of the current work is focused on traditional computing environments, with limited coverage of context-aware or resource-constrained systems. These domains introduce unique security challenges such as real-time constraints, decentralized trust assumptions, and adversarial control over physical interfaces. Although logic-based analysis has been extended to federated models and constrained networks, there is a need for logic-aware formal DSLs and compositional specification languages tailored to these environments.
@@
Scalability remains a fundamental bottleneck. While symbolic reasoning tools like SMT solvers and model checkers offer automation, their application to large-scale, real-world systems is often limited by computational overhead and manual specification burdens. Verification frameworks struggle to keep up with the dynamic, heterogeneous nature of modern cyber-infrastructures. Compositional verification—using techniques such as assume-guarantee reasoning, modular contracts, and indexed logical types—offers a potential pathway forward. However, existing tools only partially support these methods, and integration into agile development pipelines remains rare. Emerging research at the intersection of AI and formal methods offers a compelling vision for next-generation reasoning systems, but it is still in its infancy. Neural-symbolic methods for specification inference, counterexample generation, and proof synthesis show promise, yet current prototypes lack formal correctness guarantees or clear trust boundaries between learning components and deductive backends. Trust-aware pipelines, such as those based on proof-carrying code or property-carrying predictions, could help mitigate these concerns, but further theoretical development is needed.